OpenStudy (pgxgirl103):

Can you help me find the value of x? in steps would be nice so I can understand how you get it. 5x-9-16

OpenStudy (anonymous):

missing an equal sign somewheres

OpenStudy (anonymous):

if it is \[5x-9=16\] then add \(9\) and divide by \(5\)

OpenStudy (pgxgirl103):

oops, 5x-9=16 **

OpenStudy (anonymous):

k, then add 9 divide by 5

OpenStudy (pgxgirl103):

to each side?

OpenStudy (anonymous):

yes of course adding 9 gets rid of the \(-9\) on the left then dividing by \(5\) will turn \(5x\) in to \(x\) try it

OpenStudy (pgxgirl103):

so x will be 5

OpenStudy (anonymous):

yes it will lets check it by replacing \(x\) by \(5\) and see if \[5\times 5-9=16\]

OpenStudy (anonymous):

yup

OpenStudy (pgxgirl103):

thanks !